# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

### v4

@@ Discussion
One major difficulty in employing LLMs for mental health support is the difference between written and spoken communication. Hill et al. hill2015real found that people answered questions differently when asked to write the answer down when compared to when verbally expressing their answers. Future studies could include longer-term research to analyze how social penetration (i.e., the development of interpersonal relationships) over time affects information disclosure stock2023tell .
Figure 6: The challenges and potential solutions of deploying medical LLMs in clinical practice, aiming to inspire further research and development in this area.
@@ 5 Challenges
We address the challenges and discuss solutions to the adoption of LLMs in an array of medical applications (Figure 6).
@@ 5.2 Lack of Evaluation Benchmarks and Metrics
Singhal et al. medpalm proposed HealthSearchQA consisting of commonly searched health queries, offering a more human-aligned benchmark for evaluating LLM’s capabilities in the medical domain. Benchmarks such as TruthfulQA truthfulqa and HaluEval halueval evaluate more LLM-specific metrics, such as truthfulness, but do not cover the medical domain. Future research is necessary to meet the need for more medical and LLM-specific benchmarks and metrics than what are currently available.
@@ 5.3 Domain Data Limitations
Current datasets in the medical domain (Table 1) remain relatively small compared to datasets for training general-purpose LLMs (Table LABEL:table:llm_summary). These limited small datasets only cover a small space medpalm of the vase domain of medical knowledge. This results in LLMs exhibiting extraordinary performance on open benchmarks with extensive data coverage, yet falling short on real-life tasks such as differential diagnosis and personalized treatment planning medpalm2 .
@@ 5.4 New Knowledge Adaptation
LLMs are trained on extensive data to learn knowledge. Once trained, it is expensive and inefficient to inject new knowledge into an LLM through re-training. However, it is sometimes necessary to update the knowledge of the LLM, for example, on a new adverse effect of a medication or a novel disease. Two problems occur during such knowledge updates. The first problem is how to make LLMs appropriately ‘forget’ the old knowledge, as it is almost impossible to remove all ‘old knowledge’ from the training data, and the discrepancy between new and old knowledge can cause unintended association and bias hoelscher2023detecting . The second problem is the timeliness of the additional knowledge - how do we ensure the model is updated in real-time liu2023medical ? Both problems pose significant barriers to using LLMs in medical fields, where accurate and timely updates of medical knowledge are crucial in real-world implementations.
@@ 5.5 Behavior Alignment
Current solutions include instruction fine-tuning, reinforcement learning from human feedback (RLHF) instructgpt , and prompt tuning lester2021power ; liu2021p . Instruction fine-tuning zhang2023instruction refers to improving the performance of LLMs on specific tasks based on explicit instructions. For example, Ouyang et al. instructgpt used it to help LLMs generate less toxic and more suitable outputs. RLHF uses human feedback to evaluate and align the outputs of LLMs. It has been shown to be effective in multiple tasks, including becoming helpful chatbots rlhfdialogue and decision-making agents webgpt . Prompt tuning can also align LLMs to the expected output format. For example, Liu et al. chainofhindsight uses a prompting strategy, chain of hindsight, to enable the model to detect and correct its errors, thus aligning the generated output with human expectations.
5.6 Ethical, Legal and Safety Concerns
Concerns have been raised regarding using LLMs (e.g. ChatGPT) in the medical domain chatgptconcernmed , with a focus on ethics, accountability, and safety. For example, the scientific community has disapproved of using ChatGPT in writing biomedical research papers howthefuckisthisonNature due to ethical concerns. The accountability of using LLMs as assistants to practice medicine is challenging he2023survey ; tian2024opportunities . Li et al.chatgptjailbreak and Shen et al.shen2023anything found that prompt injection can cause the LLM to leak personally identifiable information (PII), e.g. email addresses, from its training data, which is a significant vulnerability when implementing LLM in the medical domain.
@@ 6 Future Directions
Although LLMs have already made a significant impact on people’s lives through chatbots and search engines, their integration into medicine is still in the infant stage. As shown in Figure 7, numerous new avenues of medical LLMs await researchers and practitioners to explore how to better serve the general public and patients.
@@ 6.1 Introduction of New Benchmarks
Recent studies have underscored the shortcomings of existing benchmarks in evaluating LLMs for clinical applications chen2023extensive ; chen2023large . Traditional benchmarks, which primarily gauge accuracy in medical question-answering, inadequately capture the full spectrum of clinical skills necessary for LLMs medpalm . Criticisms have been leveled against the use of human-centric standardized medical exams for LLM evaluation, arguing that passing these tests does not necessarily reflect an LLM’s proficiency in the nuanced expertise required in real-world clinical settings medpalm . In response, there is an emerging consensus on the need for more comprehensive benchmarks. These should include capabilities like sourcing from authoritative medical references, adapting to the evolving landscape of medical knowledge, and clearly communicating uncertainties thirunavukarasu2023large ; medpalm . Additionally, considering the sensitive nature of healthcare, these benchmarks should also assess factors such as fairness, ethics, and equity, which, though crucial, pose quantification challenges medpalm . The aim is to create benchmarks that more effectively mirror actual clinical scenarios, thus providing a more accurate measure of LLMs’ suitability for medical advisory roles.
@@ 6.2 Multimodal LLM Integrated with Time-Series, Visual, and Audio Data
Multimodal LLMs (MLLMs), or Large Multimodal Models (LMMs), are LLM-based models designed to perform multimodal (e.g. involving both visual and textual) tasks yin2023survey . While LLMs primarily address NLP tasks, MLLMs support a broader range of tasks, such as comprehending the underlying meaning of a meme and generating website codes from images. This versatility suggests promising applications of MLLMs in medicine. Several MLLM-based frameworks integrating vision and language, e.g. MedPaLM M tu2023towards , LLaVA-Med li2023llavamed , Visual Med-Alpaca Visual-Med-Alpaca , Med-Flamingo moor2023medflamingo , and Qilin-Med-VL liu2023qilin , have been proposed to adopt the medical image-text pairs for fine-tuning, thus enabling the medical LLMs to efficiently understand the input medical (e.g. radiology) images. A recent study huang2023chatgpt proposes to integrate vision, audio, and language inputs for automated diagnosis in dentistry. However, there exist only very few medical LLMs that can process time series data, such as electrocardiograms (ECGs) li2023frozen and sphygmomanometers (PPGs) englhardt2023exploring , despite such data being important for medical diagnosis and monitoring.
@@ 6.3 Medical Agents
LLM-based agents xi2023rise ; wang2023survey (i.e., intelligent systems or software designed to perform specific medical tasks or functions) have achieved significant progress in solving complex tasks (e.g. software design, molecular dynamics simulation) through human-like behaviors, such as role-playing and communication li2023camel ; tang2023medagents . However, integrating these agents effectively within the medical domain remains a challenge. The medical field involves numerous roles tang2023medagents and decision-making processes, especially in disease diagnosis that often requires a series of investigations involving CT scans, ultrasounds, electrocardiograms, and blood tests. The idea of utilizing LLMs to model each of these roles, thereby creating collaborative medical agents, presents a promising direction. These agents could mimic the roles of radiologists, cardiologists, pathologists, etc., each specializing in interpreting specific types of medical data. For example, a radiologist agent could analyze CT scans, while a pathologist agent could focus on blood test results. The collaboration among these specialized agents could lead to a more holistic and accurate diagnosis. By leveraging the comprehensive knowledge base and contextual understanding capabilities of LLMs, these agents not only interpret individual medical reports, but also integrate these interpretations to form a cohesive medical opinion.
6.4 LLMs in Medical Sub-domains
@@
Current LLM research in medicine has largely focused on general medicine, likely due to the greater availability of data in this area medpalm2 ; tian2024opportunities . This has resulted in the under-representation of LLM applications in specialized fields like ‘rehabilitation therapy’ and ‘sports medicine’. The latter, in particular, holds significant potential, given the global health challenges posed by physical inactivity. The World Health Organization identifies physical inactivity as a major risk factor for non-communicable diseases (NCDs), impacting over a quarter of the global adult population who2022physical . Despite initiatives to incorporate physical activity (PA) into healthcare systems, implementation remains challenging, particularly in developing countries with limited PA education among healthcare providers who2022physical . LLMs could play a pivotal role in these settings by disseminating accurate PA knowledge and aiding in the creation of personalized PA programs connor2023large . Such applications could significantly enhance PA levels, improving global health outcomes, especially in resource-constrained environments.
@@ 6.5 Interdisciplinary Collaborations
Just as interdisciplinary collaborations are crucial in safety-critical areas like nuclear energy production, collaborations between the medical and technology communities for developing medical LLMs are essential to ensure AI safety and efficacy in medicine. The medical community has primarily LLMs provided by technology companies without questioning their data training, which is a sub-optimal situation. Medical professionals are therefore encouraged to actively participate in creating and deploying medical LLMs by providing relevant training data, defining the desired benefits of LLMs, and conducting tests in real-world scenarios to evaluate these benefits thirunavukarasu2023large ; omiye2024large ; clusmann2023future . Such assessments would help to determine the legal and medical risks associated with LLM use in medicine and inform strategies to mitigate LLM hallucination mello2023chatgpt .
7 Conclusion
This review provides a comprehensive overview of the principles, applications, and challenges of LLMs in medicine, intended to promote further research and exploration in this interdisciplinary field. With the rapid development of foundation models, the LLMs could significantly improve future clinical practice and medical discoveries for the benefit of society. It requires sustained interdisciplinary collaboration between clinicians and AI researchers, doctor-in-the-loop, and human-centered design, to address existing challenges with applying LLMs in the medical domain. The co-development of the collaboration for appropriate training data, benchmarks, metrics, and deployment strategies could enable faster and more responsible implementation of medical large language models than the current practice.
8 Acknowledgements

### v5

@@ Discussion
Two of the most critical difficulties in employing LLMs for mental health support are the lack of emotional understanding and the risk of inappropriate or harmful responses ma2023understanding. LLMs, being language models, may struggle to fully grasp and respond to the complex emotional states and needs of individuals seeking mental health support. They may not be able to provide the same level of empathy and human connection that is crucial in therapeutic interactions.
Moreover, if not properly trained or controlled, LLMs may generate responses that are inappropriate, insensitive, or even harmful to individuals in vulnerable emotional states chung2023challenges. They may provide advice that is not grounded in evidence-based psychological practices or that goes against established mental health guidelines. Addressing these challenges requires rigorous training of LLMs in evidence-based practices, ethical considerations, and risk assessment protocols, as well as collaboration between mental health professionals and AI researchers.
4.8 Medical Inquiry and Response
The rapid advancement of LLMs also opens up new possibilities for improving healthcare delivery and patient care. LLMs, trained on vast amounts of medical knowledge, have the potential to understand and generate human-like text, making them suitable for tasks such as answering patient inquiries and assisting physicians in documentation ren2024healthcare; wang2024jmlr. As the demand for accessible and efficient healthcare services grows, researchers are exploring the use of medical LLMs to alleviate the burden on healthcare professionals and provide patients with reliable information and support. Therefore, medical inquiry and response could involve entity extraction, information retrieval, question answering, and text generation/summarization.
Guideline
Several pioneering systems demonstrate the feasibility and effectiveness of this approach. Healthcare Copilotren2024healthcare integrates dialogue, memory, and processing components to enable safe patient-LLM interactions, enhance conversations with historical data, and summarize consultations. Google’s Articulate Medical Intelligence Explore (AMIE)tu2024towards engages in diagnostic conversations with patients, exhibiting reasoning capabilities comparable to human doctors. AMIE employs a novel self-play-based simulated environment with automated feedback mechanisms to scale learning across diverse medical contexts. Another LLM-based diagnostic systemsun2024conversational enhances planning capabilities by emulating doctors, utilizing reinforcement learning for disease screening and initial diagnoses, and leveraging LLMs to parse guidelines and conduct differential diagnoses. These systems showcase the potential of medical LLMs in providing high-quality, AI-powered medical consultations and assisting physicians in their daily practice.
Discussion
However, there is still far from deploying them in the real-world healthcare system. Several challenges must be addressed before widespread deployment in real-world healthcare settings. One major concern is the potential for biased or inaccurate outputs, which could lead to improper medical advice or misdiagnosisrawte2023survey. Rigorous testing and validation across diverse patient populations and medical contexts are essential to ensure the reliability and generalizability of these systems. Additionally, the integration of medical LLMs into existing healthcare workflows and infrastructure may require substantial technical and organizational efforts. Privacy and security concerns surrounding patient data must also be carefully considered and addressed.
Furthermore, the development and deployment of medical LLMs raise important ethical and responsible AI considerations. Ensuring transparency, explainability, and accountability in the decision-making processes of these systems is crucial to maintaining trust and facilitating informed consent from patients howthefuckisthisonNature; shen2023anything. The potential impact on the doctor-patient relationship and the role of human physicians in an AI-assisted healthcare setting must also be carefully examined. Ongoing collaboration between AI researchers, healthcare professionals, ethicists, and policymakers will be necessary to establish guidelines and best practices for the responsible development and deployment of medical LLMs in real-world healthcare settings.
@@ 5 Challenges
We address the challenges and discuss solutions to the adoption of LLMs in an array of medical applications.
@@ 5.2 Lack of Evaluation Benchmarks and Metrics
Singhal et al. medpalm proposed HealthSearchQA consisting of commonly searched health queries, offering a more human-aligned benchmark for evaluating LLM’s capabilities in the medical domain. Benchmarks such as TruthfulQA truthfulqa and HaluEval halueval evaluate more LLM-specific metrics, such as truthfulness, but do not cover the medical domain. Future research is necessary to meet the need for more medical and LLM-specific benchmarks and metrics than what is currently available.
@@ 5.3 Domain Data Limitations
Current datasets in the medical domain (Table 2) remain relatively small compared to datasets for training general-purpose LLMs (Table 1). These limited small datasets only cover a small space medpalm of the vase domain of medical knowledge. This results in LLMs exhibiting extraordinary performance on open benchmarks with extensive data coverage, yet falling short on real-life tasks such as differential diagnosis and personalized treatment planning medpalm2.
@@ 5.4 New Knowledge Adaptation
LLMs are trained on extensive data to learn knowledge. Once trained, it is expensive and inefficient to inject new knowledge into an LLM through re-training. However, it is sometimes necessary to update the knowledge of the LLM, for example, on a new adverse effect of a medication or a novel disease. Two problems occur during such knowledge updates. The first problem is how to make LLMs appropriately ‘forget’ the old knowledge, as it is almost impossible to remove all ‘old knowledge’ from the training data, and the discrepancy between new and old knowledge can cause unintended association and bias hoelscher2023detecting. The second problem is the timeliness of the additional knowledge - how do we ensure the model is updated in real-time liu2023medical? Both problems pose substantial barriers to using LLMs in medical fields, where accurate and timely updates of medical knowledge are crucial in real-world implementations.
@@ 5.5 Behavior Alignment
Current solutions include instruction fine-tuning, reinforcement learning from human feedback (RLHF) instructgpt, and prompt tuning lester2021power; liu2021p. Instruction fine-tuning zhang2023instruction refers to improving the performance of LLMs on specific tasks based on explicit instructions. For example, Ouyang et al. instructgpt used it to help LLMs generate less toxic and more suitable outputs. RLHF uses human feedback to evaluate and align the outputs of LLMs. It is effective in multiple tasks, including becoming helpful chatbots rlhfdialogue and decision-making agents webgpt. Prompt tuning can also align LLMs to the expected output format. For example, Liu et al. chainofhindsight uses a prompting strategy, chain of hindsight, to enable the model to detect and correct its errors, thus aligning the generated output with human expectations.
5.6 Ethical and Safety Concerns
Concerns have been raised regarding using LLMs (e.g. ChatGPT) in the medical domain chatgptconcernmed, with a focus on ethics, accountability, and safety. For example, the scientific community has disapproved of using ChatGPT in writing biomedical research papers howthefuckisthisonNature due to ethical concerns. The accountability of using LLMs as assistants to practice medicine is challenging he2023survey; tian2024opportunities. Li et al.chatgptjailbreak and Shen et al.shen2023anything found that prompt injection can cause the LLM to leak personally identifiable information (PII), e.g. email addresses, from its training data, which is a substantial vulnerability when implementing LLM in the medical domain.
@@ 5.6 Ethical and Safety Concerns
5.7 Regulatory Challenges
The regulatory landscape of LLMs presents distinct challenges due to their large scale, broad applicability and varying reliability across applications. As LLMs progressively permeate the fields of medicine and healthcare, their versatility allows a single LLM family to facilitate a multitude of tasks across a broad spectrum of interest groups. This represents a substantial departure from the AI-based medical technologies of the past, which were typically tailored to meet specific medical needs and cater to particular interest groups newregulatoryLLM; AImed. In addition, the recent innovations of AI-enabled personalized approaches in areas such as oncology also present challenges to the traditional one-for-all auditing process newregulatoryoncology. This divergence and innovation necessitate regulators to develop adaptable, foresightful frameworks to ensure the safety, ethical standards, and privacy of the new family of LLMs-powered medical technologies.
Potential Solutions
To address the complex regulatory challenges without hindering innovation, regulators should devise adaptive, flexible, and robust frameworks. Drawing on the insights from Mesko and Topol newregulatoryLLM, creating a dedicated regulatory category and implementing patient design to enhance decision-making for LLMs used for medical purposes can better address their unique attributes and minimize harm. Furthermore, the insights outlined by Derraz et al. newregulatoryoncology emphasize the importance of implementing agile regulatory frameworks that can keep pace with the fast-paced advancements in personalized applications. Researchers both inside newregulatoryLLM; newregulatoryoncology and outside of healthcare regulatingLLM; auditingLLMs have proposed innovative strategies to regulate the use of LLMs involving (i) assessing LLMs-enabled applications in real-world settings, (ii) obligations of transparency of data and algorithms, (iii) adaptive risk assessment and mitigation processes, (iv) continuous testing and refinement of audited technologies. Such proactive regulatory adaptations are crucial to maintaining high standards of safety, ethics, and trustworthiness of medical technology.
@@ 6 Future Directions
Although LLMs have already made an impact on people’s lives through chatbots and search engines, their integration into medicine is still in the infant stage. As shown in Figure 6, numerous new avenues of medical LLMs await researchers and practitioners to explore how to better serve the general public and patients.
@@ 6.1 Introduction of New Benchmarks
Recent studies have underscored the shortcomings of existing benchmarks in evaluating LLMs for clinical applications chen2023extensive; chen2023large. Traditional benchmarks, which primarily gauge accuracy in medical question-answering, inadequately capture the full spectrum of clinical skills necessary for LLMs medpalm. Criticisms have been leveled against the use of human-centric standardized medical exams for LLM evaluation, arguing that passing these tests does not necessarily reflect an LLM’s proficiency in the nuanced expertise required in real-world clinical settings medpalm. In response, there is an emerging consensus on the need for more comprehensive benchmarks. These should include capabilities like sourcing from authoritative medical references, adapting to the evolving landscape of medical knowledge, and clearly communicating uncertainties thirunavukarasu2023large; medpalm. To further enhance the relevance of these benchmarks, new benchmarks should incorporate scenarios that test an LLM’s ability through simulation of real-world applications and adjust to feedback from clinicians while maintaining robustness. Additionally, considering the sensitive nature of healthcare, these benchmarks should also assess factors such as fairness, ethics, and equity, which, though crucial, pose quantification challenges medpalm. While efforts such as the AMIE study have advanced benchmarking by utilizing real physician evaluations and comprehensive criteria rooted in actual clinical skills and communication, as reflected in the Objective Structured Clinical Examination (OSCE), there remains a pressing need for benchmarks that are adaptive, scalable and robust for other diverse and personalized applications of LLMs. The aim is to create benchmarks that more effectively mirror diverse real-world clinical scenarios, thus providing a more accurate measure of LLMs’ suitability for their applications in medicine. Future research may focus on (i) using synthetic data along with real-world data to create benchmarks that are both comprehensive and scalable, (ii) using clinical guidelines and criteria to reflect real-world values that are not normally included in traditional benchmarks, (iii) physician-in-the-loop benchmarks to evaluate the performance of LLMs leveraging their human counterparts or users.
@@ 6.2 Multimodal LLM Integrated with Time-Series, Visual, and Audio Data
Multimodal LLMs (MLLMs), or Large Multimodal Models (LMMs), are LLM-based models designed to perform multimodal (e.g. involving both visual and textual) tasks yin2023survey. While LLMs primarily address NLP tasks, MLLMs support a broader range of tasks, such as comprehending the underlying meaning of a meme and generating website codes from images. This versatility suggests promising applications of MLLMs in medicine. Several MLLM-based frameworks integrating vision and language, e.g. MedPaLM M tu2023towards, LLaVA-Med li2023llavamed, Visual Med-Alpaca Visual-Med-Alpaca, Med-Flamingo moor2023medflamingo, and Qilin-Med-VL liu2023qilin, have been proposed to adopt the medical image-text pairs for fine-tuning, thus enabling the medical LLMs to efficiently understand the input medical (e.g. radiology) images. A recent study huang2023chatgpt proposes to integrate vision, audio, and language inputs for automated diagnosis in dentistry. However, there exist only very few medical LLMs that can process time series data, such as electrocardiograms (ECGs) li2023frozen and sphygmomanometers (PPGs) englhardt2023exploring, despite such data being important for medical diagnosis and monitoring. Although early in their proposed research stages, these studies suggest that MLLMs trained at scale have the potential to effectively generalize across various domains and modalities outside of NLP tasks. However, the training of MLLMs at scale is still costly and ineffective, resulting in the size of MLLMs being much smaller than LLMs. Moving forward, future research may focus on (i) more effective processing, representation, and learning of multi-modal data and knowledge, (ii) cost-effective training of MLLMs, especially modalities that are more resource-demanding such as videos and images, (iii) collecting or accessing safely, currently unavailable, multi-modal data in medicine and healthcare.
@@ 6.3 Medical Agents
LLM-based agents xi2023rise; wang2023survey utilize LLMs as controllers to leverage their reasoning capabilities. By integrating LLMs with external tools and multimodal perceptions, these agents can interact with environments, learn from feedback, and acquire new skills, enabling them to solve complex tasks (e.g., software design, molecular dynamics simulation) through human-like behaviors, such as role-playing and communication li2023camel; tang2023medagents.
@@
However, integrating these agents effectively within the medical domain remains a challenge. The medical field involves numerous roles tang2023medagents and decision-making processes, especially in disease diagnosis that often requires a series of investigations involving CT scans, ultrasounds, electrocardiograms, and blood tests. The idea of utilizing LLMs to model each of these roles, thereby creating collaborative medical agents, presents a promising direction. These agents could mimic the roles of radiologists, cardiologists, pathologists, etc., each specializing in interpreting specific types of medical data. For example, a radiologist agent could analyze CT scans, while a pathologist agent could focus on blood test results. The collaboration among these specialized agents could lead to a more holistic and accurate diagnosis. By leveraging the comprehensive knowledge base and contextual understanding capabilities of LLMs, these agents not only interpret individual medical reports but also integrate these interpretations to form a cohesive medical opinion. To enhance the integration of LLMs-based agents, future research may explore (i) a seamless data pipeline that collects data from various devices and transforms them into data format compatible with LLMs (ii) effective communication and collaboration between agents, especially in areas such as ensuring truthfulness during communication, dispute resolution between agents, and role-based data security measures, (iii) real-time decision-making such as making timely decisions using data collected from remote monitoring devices, (iv) adaptive learning such as preparing for a new pandemic or learning from unseen medical conditions.
6.4 LLMs in Underrepresented Specialties
@@
Current LLM research in medicine has largely focused on general medicine, likely due to the greater availability of data in this area medpalm2; tian2024opportunities. This has resulted in the under-representation of LLM applications in specialized fields like ‘rehabilitation therapy’ or ‘sports medicine’. The latter, in particular, holds potential, given the global health challenges posed by physical inactivity. The World Health Organization identifies physical inactivity as a major risk factor for non-communicable diseases (NCDs), impacting over a quarter of the global adult population who2022physical. Despite initiatives to incorporate physical activity (PA) into healthcare systems, implementation remains challenging, particularly in developing countries with limited PA education among healthcare providers who2022physical. LLMs could play a pivotal role in these settings by disseminating accurate PA knowledge and aiding in the creation of personalized PA programs connor2023large. Such applications could enhance PA levels, improving global health outcomes, especially in resource-constrained environments. To spark innovation in these underrepresented specialties, future research can focus on areas such as (i) effective data collection in underrepresented specialties, (ii) applications of LLMs in assisting with tasks of underrepresented specialties, (iii) using LLMs to help progress the research of these underrepresented specialties.
@@ 6.5 Interdisciplinary Collaborations
Just as interdisciplinary collaborations are crucial in safety-critical areas like nuclear energy production, collaborations between the medical and technology communities for developing medical LLMs are essential to ensure AI safety and efficacy in medicine. The medical community has primarily adopted LLMs provided by technology companies without rigorously questioning their data training, ethical protocols, or privacy protection. Medical professionals are therefore encouraged to actively participate in creating and deploying medical LLMs by providing relevant training data, defining the desired benefits of LLMs, and conducting tests in real-world scenarios to evaluate these benefits thirunavukarasu2023large; omiye2024large; clusmann2023future. Such assessments would help to determine the legal and medical risks associated with LLM use in medicine and inform strategies to mitigate LLM hallucination mello2023chatgpt. Additionally, training ‘bilingual’ professionals—those versed in both medicine and LLM technology—is increasingly vital due to the rapid integration of LLMs in healthcare. Future research may explore (i) interdisciplinary frameworks, such as frameworks to facilitate the sharing of localized data from rural clinics, (ii) ‘bilingual education programs’ that offer training from both worlds - AI and medicine, (iii) effective in-house development methods to help hospitals and physicians ‘guard’ patient data from corporations while still being able to embrace innovation.
Acknowledgements